# Optimizing Posterior Samples for Bayesian Optimization via Rootfinding

### v1

@@ 6 Conclusion and Future Work
We presented TS-roots, a global optimization strategy for posterior sample paths. It comprises an adaptive selection of starting points for gradient-based multistart optimizers, combining exploration with exploitation. This strategy breaks the curse of dimensionality by exploiting the separability of Gaussian process priors. Compared to random multistart and a genetic algorithm, TS-roots consistently yields higher-quality solutions in optimizing posterior sample-based acquisition functions for Bayesian optimization, both in low- and high-dimensional settings. It also improves the outer-loop optimization performance of GP-TS and information-theoretic acquisition functions such as MES. For future work, we extend TS-roots to other spectral representations per Bochner’s theorem [16, 35, 36]. We also plan to study the modes and the probability when TS-roots fails to find the global optimum, and the impact of subset sizes.
@@ A.1 Proof of Proposition 1: A Representation of the Set of Local Minima
For a boundary point $\mathbf{x}\in\partial\mathcal{X}:=\mathcal{X}\setminus\operatorname*{int}\mathcal{X}$, we partition the index set $\{1,\cdots,d\}$ into $L,R$, and $I$ such that $x_{i}=\underline{x}_{i}$ for all $i\in L$, $x_{i}=\overline{x}_{i}$ for all $i\in R$, and $x_{i}\in(\underline{x}_{i},\overline{x}_{i})$ for all $i\in I$. Define $\nabla_{J}:=(\partial_{j})_{j\in J}$ for any subset $J$ of the indices. Then $\mathbf{x}$ is a strong local minimum of $f$ if and only if the following conditions hold: (a) $\mathbf{x}$is a strong local minimum in $\{x_{j}\}_{j\notin I}\times\prod_{i\in I}[\underline{x}_{j},\overline{x}_{j}]$; (b) $\nabla_{L}f(\mathbf{x})>0$; and (c) $\nabla_{R}f(\mathbf{x})<0$.
@@ A.2 Number of Local Minima of a Separable Function
In Proposition 1, each set of candidate coordinates $\Xi_{i}$ is partitioned into mono type and mixed type:
@@ C.1 Spectral Sampling of Separable Gaussian Process Priors
Given spectral representations of the univariate component functions of a separable Gaussian Process prior, we can accurately approximate the prior sample as
@@
Here $N_{i}$ is selected for each variate such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}$ is sufficiently small (see Appendix F for the value used in this study). Using spectral representations of the univariate components as in Eq. 10 is much more efficient than directly using a spectral representation of the separable GP prior, because the former uses $\sum_{i=1}^{d}N_{i}$ univariate terms to exactly represent $\prod_{i=1}^{d}N_{i}$ multivariate terms in the latter.
@@
The univariate squared exponential (SE) covariance function can be written as $\kappa(x,x^{\prime};l)=\exp(-\frac{1}{2}s^{2})$, where the relative distance $s=|x-x^{\prime}|/l$. The spectral representation of such a covariance function per Mercer’s theorem is $\kappa(x,x^{\prime})=\sum_{k=0}^{\infty}\lambda_{k}\phi_{k}(x)\phi_{k}(x^{\prime})$. With a Gaussian measure $\mu=\mathcal{N}(0,\sigma^{2})$ over the domain $\mathcal{X}=\mathbb{R}$, we can write the eigenvalues $\lambda_{k}$ and eigenfunctions $\phi_{k}(x)$ of the kernel integral operator as follows. (See e.g., [45] Sec. 4 and [46] 7.374 eq. 8.)
@@ C.2 Univariate Global Rootfinding
Algorithm 2 roots: Univariate global rootfinding on an interval.
@@
A proof of Theorem 1 is provided in [33], Chapter 18. A classical formula to compute the weights $\{a_{k}\}$ requires $\mathcal{O}{(m^{2})}$ floating point operations, which can be reduced to $\mathcal{O}{(m\log m)}$ using a fast Fourier transform. Since the colleague matrix is tridiagonal except in the final row, the complexity of computing its eigenvalues can be improved from $\mathcal{O}{(m^{3})}$ to $\mathcal{O}{(m^{2})}$ operations, which can be further improved to $\mathcal{O}{(m)}$ via recursive subdivision of intervals (see Trefethen [33]). The roots algorithm is implemented in the Chebfun package in MATLAB [47] and the chebpy package in Python [48]; both packages also implement other related programs such as chebfun for Chebyshev polynomial approximation and diff for differentiation.
@@ C.3 Best Local Minima of a Separable Function
Given the univariate component functions of a separable function, Algorithm 3 finds the subset $S_{\text{o}}$ of the local minima of the function with the $n_{o}$ smallest function values. This procedure requires the maxk_sum algorithm in Algorithm 1, the roots algorithm in Algorithm 2 and the related programs chebfun and diff, see also Appendix F.
@@ C.4 Decoupled Sampling from Gaussian Process Posteriors
Algorithm 7 Gaussian process Thompson sampling acquisition function
1:Gaussian process posterior $f^{k}$
2:$\alpha^{k}(\mathbf{x})\sim f^{k}$
Appendix D Bayesian Optimization via Thompson Sampling
@@
Appendix E Benchmark Functions
@@
Appendix F Experimental Details
@@
We use squared exponential (SE) covariance functions for our experiments. The spectra of univariate SE covariance functions for all problems (see Section C.1) are determined using the Gaussian measure $\mu=\mathcal{N}(0,1)$. The number of terms $N_{i}$, $i\in\{1,\cdots,d\}$, of each truncated univariate spectrum is determined such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}=10^{-16}$. If $N_{i}>1000$, we set $N_{i}=1000$ to trade off between the accuracy of truncated spectra and computational cost. The maximum size of the exploration set is $n_{\text{e}}=250$. The maximum size of the exploitation set is $n_{\text{x}}=200$.
@@
Appendix G Additional Results
@@
Figure 8 shows how we can improve the exploitation of GP-TS when increasing the exploration-exploitation control parameter $N_{\text{c}}$.
@@
Figure 9 shows the performance of sample-average TS-roots with different exploration-exploitation control parameters $N_{\text{c}}$ for the 2D Schwefel, 4D Rosenbrock, and 6d Ackley functions.

### v2

@@ 7 Conclusion and Future Work
We presented TS-roots, a global optimization strategy for posterior sample paths. It features an adaptive selection of starting points for gradient-based multistart optimizers, combining exploration and exploitation. This strategy breaks the curse of dimensionality by exploiting the separability of Gaussian process priors. Compared with random multistart and a genetic algorithm, TS-roots consistently yields higher-quality solutions in optimizing posterior sample paths, across a range of input dimensions. It also improves the outer-loop optimization performance of GP-TS and information-theoretic acquisition functions such as MES for Bayesian optimization. For future work, we aim to extend TS-roots to other spectral representations per Bochner’s theorem [17, 39, 40]. We also plan to study the ways and the probability of TS-roots failing to find the global optimum, as well as the impact of subset sizes.
@@ A.1 Proof of Proposition 1: A Representation of the Set of Local Minima
For a boundary point $\mathbf{x}\in\partial\mathcal{X}:=\mathcal{X}\setminus\operatorname*{int}\mathcal{X}$, we partition the index set $\{1,\cdots,d\}$ into $L,R$, and $I$ such that $x_{i}=\underline{x}_{i}$ for all $i\in L$, $x_{i}=\overline{x}_{i}$ for all $i\in R$, and $x_{i}\in(\underline{x}_{i},\overline{x}_{i})$ for all $i\in I$. Define $\nabla_{J}:=(\partial_{j})_{j\in J}$ for any subset $J$ of the indices. Then $\mathbf{x}$ is a strong local minimum of $f$ if and only if the following conditions hold: (a) $\mathbf{x}$is a strong local minimum in $\{x_{j}\}_{j\notin I}\times\prod_{j\in I}[\underline{x}_{j},\overline{x}_{j}]$; (b) $\nabla_{L}f(\mathbf{x})>0$; and (c) $\nabla_{R}f(\mathbf{x})<0$.
@@ A.2 Number of Local Minima of a Separable Function
In Proposition 1, each set of candidate coordinates $\Xi_{i}$ is partitioned into mixed type and mono type:
@@ C.1 Spectral Sampling of Separable Gaussian Process Priors
Given spectral representations of the univariate component functions of a separable Gaussian Process prior, we can accurately approximate the prior sample as:
@@
Here $N_{i}$ is selected for each variate such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}$ is sufficiently small (see Appendix H for the value used in this study). Using spectral representations of the univariate components as in Eq. 10 is much more efficient than directly using a spectral representation of the separable GP prior, because the former uses $\sum_{i=1}^{d}N_{i}$ univariate terms to exactly represent $\prod_{i=1}^{d}N_{i}$ multivariate terms in the latter.
@@
The univariate squared exponential (SE) covariance function can be written as $\kappa(x,x^{\prime};l)=\exp(-\frac{1}{2}s^{2})$, where the relative distance $s=|x-x^{\prime}|/l$ and length scale $l\in(0,\infty)$. The spectral representation of such a covariance function per Mercer’s theorem is $\kappa(x,x^{\prime})=\sum_{k=0}^{\infty}\lambda_{k}\phi_{k}(x)\phi_{k}(x^{\prime})$. With a Gaussian measure $\mu=\mathcal{N}(0,\sigma^{2})$ over the domain $\mathcal{X}=\mathbb{R}$, we can write the eigenvalues $\lambda_{k}$ and eigenfunctions $\phi_{k}(x)$ of the kernel integral operator as follows. (See e.g., [48] Sec. 4 and [49] 7.374 eq. 8.)
@@ C.2 Univariate Global Rootfinding
Algorithm 3 roots: Univariate global rootfinding on an interval
@@
A proof of Theorem 1 is provided in [37], Chapter 18. A classical formula to compute the weights $\{a_{k}\}$ requires $\mathcal{O}{(m^{2})}$ floating point operations, which can be reduced to $\mathcal{O}{(m\log m)}$ using a fast Fourier transform. Since the colleague matrix is tridiagonal except in the final row, the complexity of computing its eigenvalues can be improved from $\mathcal{O}{(m^{3})}$ to $\mathcal{O}{(m^{2})}$ operations, which can be further improved to $\mathcal{O}{(m)}$ via recursive subdivision of intervals (see Trefethen [37]). Specifically, if $m>100$, the interval is divided recursively so that on each subinterval the function can be accurately approximated by a polynomial of degree no greater than 100. The roots algorithm is implemented in the Chebfun package in MATLAB [50] and the chebpy package in Python [51]; both packages also implement other related programs such as chebfun for Chebyshev polynomial approximation and diff for differentiation.
@@ C.3 Best Local Minima of a Separable Function
Given the univariate component functions of a separable function, Algorithm 4 finds the subset $S_{\text{o}}$ of the local minima of the function with the $n_{o}$ smallest function values. This procedure requires the maxk_sum algorithm in Algorithm 2, the roots algorithm in Algorithm 3 and the related programs chebfun and diff, see also Appendix H.
@@ C.4 Decoupled Sampling from Gaussian Process Posteriors
C.5 Computational Complexity of TS-roots
Per Algorithm 1, the computational cost of the TS-roots method is dominated by a few tasks: (1) one call of minsort (Algorithm 4); (2) $n_{\text{o}}+n$ evaluations of the posterior sample path $\widetilde{f}(\cdot)$; and (3) $n_{\text{e}}+n_{\text{x}}$ calls of the gradient-based optimizer minimize.
First, consider task (2). Evaluating $\widetilde{f}(\cdot)$ involves evaluating: (i) the prior sample path $f(\cdot)$, which involves evaluating its $d$ univariate component functions, each with a cost that depends on its spectral representation (Section C.1); and (ii) the canonical basis $\boldsymbol{\kappa}_{\cdot,n}(\cdot)$, which costs $\mathcal{O}(dn)$ flops. When the data size $n$ is large, we can pre-filter the observed locations $X$ by the observations $\mathbf{y}$, which is a good estimate of $\widetilde{f}(X)$ depending on the observation noise. We assume that the number of observed locations after filtering is at most comparable to $n_{\text{e}}$. The cost of task (2) is thus $\mathcal{O}(n_{\text{o}}dn)$ flops.
Now consider task (3). Evaluating the gradient of $\widetilde{f}(\cdot)$ involves evaluating the gradients of $f(\cdot)$ and $\boldsymbol{\kappa}_{\cdot,n}(\cdot)$. Since both $f(\cdot)$ and $\boldsymbol{\kappa}_{\cdot,n}(\cdot)$ are separable functions, their gradients can be computed at a cost comparable to evaluating their function values. Let $N_{\text{grad}}$ be the number of gradient evaluations required by the gradient-based optimizer. The cost of task (3) is thus $\mathcal{O}((n_{\text{e}}+n_{\text{x}})N_{\text{grad}}dn)$ flops.
For task (1), the cost of the minsort algorithm is dominated by: (i) $d$ calls to chebfun, which evaluates the univariate components $f_{i}(\cdot)$ at a number of points depending on their complexity; (ii) $d$ calls to roots (Algorithm 3), which scales linearly with the polynomial degree $m$ of the chebfun object, itself dependent on the complexity of $f_{i}(\cdot)$; and (iii) at most one call to maxk_sum (Algorithm 2) at a cost of $\mathcal{O}(tn_{\text{o}}\log n_{\text{o}})$, where $t=\sum_{i=1}^{d}t_{i}$ and $t_{i}$ is two plus the number of critical points of $f_{i}(\cdot)$ which depends on the complexity of $f_{i}(\cdot)$. The complexity of $f_{i}(\cdot)$, for the SE kernel for example (see Section C.1), can be quantified as the inverse length scale $\theta_{i}=1/l_{i}$. We may define an average complexity as $\theta=\frac{1}{d}\sum_{i=1}^{d}\theta_{i}$. The cost of task (1) is thus $\mathcal{O}(d\theta n_{\text{o}}\log n_{\text{o}})$ flops.
As explained in Appendix D, we can set $n_{\text{e}}$ and $n_{\text{x}}$ to small values and $n_{\text{o}}$ to a moderate value, independent of $\widetilde{f}(\cdot)$ and thus independent of $d$, $n$ and $\theta$. The overall cost of TS-roots thus scales as $\mathcal{O}(dn+d\theta)$, which is linear in the input dimension $d$.
Appendix D Minimum Size of Exploration and Exploitation Sets
We conduct an empirical experiment to determine minimal values for $n_{\text{e}}$ and $n_{\text{x}}$ of TS-roots algorithm, which are the sizes of $S_{\text{e}}$ and $S_{\text{x}}$, respectively. From this experiment, we also recommend a value for $n_{\text{o}}$, which is the size of $S_{\text{o}}$. Recall that points in $S_{\text{o}}$ are sorted in ascending order of prior sample values, while those in $S_{\text{e}}$ and $S_{\text{x}}$ are sorted in ascending order of posterior sample values.
Let $I_{\text{e}}$ and $I_{\text{x}}$ be the sets of indices of points in $S_{\text{e}}$ and $S_{\text{x}}$ that converge to the best local minimum of the posterior sample in each optimization iteration, respectively. Let $I_{\text{o}}$ be the set of indices of points in $S_{\text{o}}$ associated with $S_{\text{e}}$. Our hypothesis is that we have a high chance of finding a small index value in either $I_{\text{e}}$ or $I_{\text{x}}$. If this hypothesis is confirmed, then we can set both $n_{\text{e}}$ and $n_{\text{x}}$ at very small values, which significantly accelerate the inner-loop optimization. To confirm our hypothesis, we employ the following two steps. First, we set $n_{\text{e}}$ and $n_{\text{x}}$ at large values to mimic the effect of removing the set size limits, ensuring accurate solution of the global optimization problem. Then, we show that we have a high chance of finding a small index value from $I_{\text{e}}$ and/or $I_{\text{x}}$ in each optimization iteration.
Figure 7: Left column: Minimum index values and index range of $I_{\text{e}}$ and/or $I_{\text{x}}$ for starting points that converge to the best local minimum $\bf{x}^{\star}$ of posterior sample in each optimization iteration, and index values of $I_{\text{o}}$ associated with minimum index values from $I_{\text{e}}$. Middle column: Zoom-in plots of index values. Right column: Historam of the minimum index values. (a) 2D Schwefel, (b) 4D Rosenbrock, (c) 10D Levy, (d) 16D Ackley, (e) 16D Powell functions.
We test our hypothesis on the 2D Schwefel, 4D Rosenbrock, 10D Levy, 16D Ackley, 16D Powell functions. We set $n_{\text{o}}=5000$, $n_{\text{e}}=n_{\text{x}}=1000$, and $\alpha=3$ (buffer coefficient). The left and middle columns of Figure 7 show the smallest index values and the variation of index values from $I_{\text{e}}$ and/or $I_{\text{x}}$ of starting points that converge to the best local minimum $\bf{x}^{\star}$ of the posterior sample path in each optimization iteration. The left column also plots the index values from $I_{\text{o}}$ corresponding to the smallest index values from $I_{\text{e}}$, if exists. The right column shows the histograms of the smallest index values from $I_{\text{e}}$ and $I_{\text{x}}$ for all iterations considered. These results show that we have a high chance of finding a small index value from $I_{\text{e}}$ and/or $I_{\text{x}}$ in each iteration. This confirms our hypothesis. In fact, using the first point in $S_{\text{e}}$ and the first point in $S_{\text{e}}$—only two points—we can discover the global optimum most of the time. Interestingly, the smallest index values appear largely independent of both the optimization iteration and the input dimension. Furthermore, the results suggest that it is safe to set $n_{\text{o}}=500$; and for almost exact global optimization, we suggest setting $n_{\text{e}}=25$ and $n_{\text{x}}=50$.
@@ Appendix D Minimum Size of Exploration and Exploitation Sets
Appendix E Bayesian Optimization via Thompson Sampling
@@
Appendix F Benchmark Functions
@@
Figure 8: Ten-bar truss. Cross-sectional areas of ten truss members are the input variables $x_{i}$, $i\in\{1\dots,10$}. Known parameters include length $L$, Young’s modulus of truss material $E$, and external loads $P_{j}$, $j\in\{1,2,3\}$. The vertical displacement at node 3 is denoted as $\delta_{3}$.
Appendix G Ten-bar Truss
Consider a ten-bar truss shown in Figure 8. The truss has ten members and is subjected to vertical load $P_{1}=60$ kN at node 2, vertical load $P_{2}=40$ kN at node 3, and horizontal load $P_{3}=40$ kN at node 3. The Young’s modulus of the truss material $E=200$ GPa. The length parameter $L=1$ m. Let $A({\bf x})=\sum_{i=1}^{10}x_{i}$ and $\delta_{3}({\bf x})$ denote the total area of the cross-sectional areas of the truss members and the vertical displacement at node 3, respectively, where ${\bf x}=[x_{1},\dots,x_{10}]^{\intercal}$ is the vector of cross-sectional areas of the truss members. The optimization problem formulated for the truss is to minimize both $A({\bf x})$ and $\delta_{3}({\bf x})$. Since $A({\bf x})$ and $\delta_{3}({\bf x})$ are competing, we define the objective function as a weight-sum of $A({\bf x})$ and $\delta_{3}({\bf x})$, such that
$f({\bf x})=w_{1}\frac{A({\bf x})}{A_{\max}}+w_{2}\frac{\delta_{3}({\bf x})}{\delta_{\max}},$ (19)
where ${\bf x}\in[1,20]^{10}\,\mathrm{cm}^{2}$, $w_{1}=0.6$, $w_{2}=0.4$, $A_{\max}=200\,\mathrm{cm}^{2}$, and $\delta_{\max}=3\,\mathrm{cm}$.
Appendix H Experimental Details
@@
We use squared exponential (SE) covariance functions for our experiments. The spectra of univariate SE covariance functions for all problems (see Section C.1) are determined using the Gaussian measure $\mu=\mathcal{N}(0,1)$. The number of terms $N_{i}$, $i\in\{1,\cdots,d\}$, of each truncated univariate spectrum is determined such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}=10^{-16}$. If $N_{i}>1000$, we set $N_{i}=1000$ to trade off between the accuracy of truncated spectra and computational cost. We also set $n_{\text{o}}=500$. The maximum size of the exploration set is $n_{\text{e}}=250$. The maximum size of the exploitation set is $n_{\text{x}}=200$.
@@
Appendix I Additional Results
@@
Comparison of Inner-loop Optimization Results.
Figures 10 and 11 compare the performance of the inner-loop optimization by three different initialization schemes, i.e., rootfinding, uniform grid, and Latin hypercube sampling, for low-dimensional cases of the 2D Schwefel and 4D Rosenbrock functions, and for higher-dimensional cases of the 10D Levy, 16D Ackley, and 16D Powell functions. Rootfinding performs better than the uniform grid and Latin hypercube sampling initialization schemes, especially in high-dimensional settings.
@@
Figure 12 shows how we can improve the exploitation of GP-TS when increasing the exploration–exploitation control parameter $N_{\text{c}}$.
@@
Figure 13 shows the performance of sample-average TS-roots with different exploration–exploitation control parameters $N_{\text{c}}$ for the 2D Schwefel, 4D Rosenbrock, and 6D Ackley functions.
@@
Figure 10: Inner-loop optimization results by three different initialization schemes, i.e., rootfinding, uniform grid, and Latin hypercube sampling, for (a) the 2D Schwefel and (b) 4D Rosenbrock functions. The plots are cumulative values of optimized GP-TS acquisition functions $\alpha_{k}^{\star}$, cumulative distances between new solution points $\mathbf{x}_{k}^{\star}$ and the true global minima $\mathbf{x}_{k}^{\text{t}}$ of the acquisition functions, and cumulative CPU times $t_{k}$ for optimizing the acquisition functions.
Figure 11: Inner-loop optimization results by three different initialization schemes, i.e., rootfinding, uniform grid, and Latin hypercube sampling, for (a) the 10D Levy, (b) 16D Ackley, and (c) 16D Powell functions. The plots are cumulative values of optimized GP-TS acquisition functions $\alpha_{k}^{\star}$ and cumulative CPU times $t_{k}$ for optimizing the acquisition functions.